OpenStudy (anonymous):

If sin (B) = -1/3 with B in QIII, find csc (B/2).

OpenStudy (anonymous):

\[\sin \left( \frac{B}{2} \right)=\sqrt{\frac{1-\cos B}{2}}\]

OpenStudy (anonymous):

Right, but how does this help find csc?

OpenStudy (anonymous):

\[\csc \left( \frac{B}{2} \right)=\frac{1}{\sin \left( \frac{B}{2} \right)}\]

OpenStudy (anonymous):

Oh!!! I didn't think about it that way... So, if it's B/2, would I do (-1/3)/2?

OpenStudy (anonymous):

no, it doesnt work that way, dividing the angle by 2 is not the same as dividing the value by 2

OpenStudy (anonymous):

Good point. I'm sorry. This stuff is so confusing to me.

OpenStudy (anonymous):

given sin B = -1/3, you need to find cos B first, and then plug it in into the half angle formula

OpenStudy (anonymous):

Ok... So, to find cos B, I use the formula\[\pm \sqrt{(1+cosB)/2}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):